FXY equals six, YZ equals four and XZ equals to which point is between the other two
1
Expert's answer
2020-12-17T18:52:20-0500

XY = 6 YZ = 4 XZ = 2


XY = XZ + YZ "\implies" 6 = 2 + 4 "\implies" 6 = 6


So ,


Midpoint is Z because XY = XZ+YZ